OpenStudy (anonymous):

Identify the 10th term of the geometric sequence 7, –35, 175 ? can someon show me how to do this please

OpenStudy (asnaseer):

ok - first you need to work out the common ratio between each term. so, what do you need to multiply 7 by in order to get "-35"?

OpenStudy (anonymous):

-5

OpenStudy (asnaseer):

correct. now as a quick check you can confirm that: (-35) * (-5) = 175

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so now we know the common ratio and we also know the first term of the sequence commonly referred to as 'a' is equal to 7. in summary, we now know: a=7 r=-5

OpenStudy (anonymous):

okayy

OpenStudy (asnaseer):

next, you need to use the formulae for the n'th term of a geometric sequence. do you know the formula?

OpenStudy (anonymous):

no i dont,

OpenStudy (asnaseer):

it might help you if you first refer to a quick refresher on geometric sequences. here is one: http://www.mathsisfun.com/algebra/sequences-sums-geometric.html once you are ready, I can help you further if you still require help.

OpenStudy (anonymous):

so do you just muliply each term by -5 ?

OpenStudy (asnaseer):

correct. the /long/ way of doing this would be to keep multiplying by -5 until you get to the 10'th term. but you can use the formula for the n'th term and use that directly to get the 10'th term.

OpenStudy (anonymous):

can you tell me the formula ? or is it the same formula for a arthmetic sequence ?

OpenStudy (asnaseer):

it is not the same as arithmetic sequences. the formula is given on that page if you scroll down a bit and look at the section titled "The Rule"

OpenStudy (asnaseer):

\[x_n=ar^{n-1}\]

OpenStudy (anonymous):

Xn= ar ^(n-1) ?

OpenStudy (asnaseer):

yes - here 'a' is the first term r is the common ratio n is the index of the term you want - in this case 10

OpenStudy (anonymous):

why 10 ?

OpenStudy (asnaseer):

because the question is asking you to find the 10'th term of the sequence

OpenStudy (anonymous):

oh okay

OpenStudy (asnaseer):

so using the numbers we found, we get:\[x_{10}=7*(-5)^{10-1}=7*(-5)^9\]

OpenStudy (asnaseer):

and since 9 is an odd power, we know we must end up with a negative number. therefore:\[x_{10}=-7*5^9\]which you should be able to calculate using a calculator.

OpenStudy (anonymous):

13,671,875 ?

OpenStudy (asnaseer):

remember the negative sign

OpenStudy (anonymous):

-13,671,875 *

OpenStudy (asnaseer):

thats it - you got it! - well done!

OpenStudy (anonymous):

so if the exponent is odd the answer will be negative ?

OpenStudy (asnaseer):

only if the common ratio is negative - which it was in this case (-5)

OpenStudy (anonymous):

so if the common ratio is negative the answer is negative ?

OpenStudy (asnaseer):

if the common ration is negative, then the terms alternate between positive and negative values. you can see that from the first 3 terms given to you: 7, -35, 175

OpenStudy (anonymous):

thanks you

OpenStudy (asnaseer):

yw